Lupe made a down payment of $1500 toward the purchase of a new car. To pay the balance of the purchase price, she has secured a loan from her bank at the rate of 10%/year compounded monthly. Under the terms of her finance agreement she is required to make payments of $220/month for 36 months. What was the cash price of the car? (Round your answer to the nearest cent.)
$
Answer #1

Loan amount = Present value of Annuity = A*[(1-(1+r)-n)/r]

Where

A - Annuity payment = 220

r - rate per period = .10/12

n - no. of periods = 36

Loan amount = 220*[(1-(1+(.10/12))^-36)/(.10/12)]

= 220*30.9912355853

= 6818.07

cash price of the car = Loan amount +down payment

= 6818.07+1500

= 8318.07